OpenStudy (anonymous):

find specified scalar: u=-8i+5j and v=-11i-6j; find u*v

jimthompson5910 (jim_thompson5910):

Dot Product if u = ai + bj and v = ci + dj then u dot v = a*c + b*d

jimthompson5910 (jim_thompson5910):

Example: u = 2i - 3j v = 7i + 5j u dot v = 2*7 + (-3)*5 u dot v = 14 - 15 u dot v = -1

OpenStudy (anonymous):

right =(-8)(-11) + 5(-6) which equals -62; but what is the specified scalar?

jimthompson5910 (jim_thompson5910):

(-8)(-11) + 5(-6) = 58, not -62

jimthompson5910 (jim_thompson5910):

the specified scalar would be 58

jimthompson5910 (jim_thompson5910):

scalar is just a single number (ie not a vector)

OpenStudy (anonymous):

AH!! Thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome